Daily Report - Yosemite National Park
Tuesday, August 15th, 2017
Today: Sunny. Near the crest, a 20 percent chance of showers in the afternoon. Highs 74 to 79 at 5000 feet...65 to 70 at 8000 feet.
Tonight: Partly cloudy. Near the crest, a 20 percent chance of showers in the evening. Lows 54 to 59 at 5000 feet...47 to 52 at 8000 feet.
Tomorrow: Mostly cloudy. Slight chance of showers in the afternoon. Near the crest, a 30 percent chance of showers in the afternoon. Highs 77 to 82 at 5000 feet...68 to 73 at 8000 feet.
Additional Point Forecast Weather Links:
NEW AND HAPPENING TODAY
Annual Big Time Event This Weekend
The Annual Big Time Event will be at the Village of Ahwahnee behind the Valley Museum from August 18 through August 20. Please remove all vehicles from the parking area adjacent to the Emergency Services Building, gas pumps, and AT&T building from the evening of Friday August 18 through Sunday August 20. Because of the limited parking due to the on going construction at the Emergency Services Building, employees can free up event parking on Village Drive by parking up near the NPS Stables or at the courthouse on Saturday and Sunday. (J. Hoeflich)
Road Work this Friday night at "The Narrows"
Contractors need to work one Friday night (8/18) this month at “the Narrows” (El Portal Road between Pohono Bridge and the Big Oak Flat intersection) to pour concrete. Delays will be:
8pm to 10pm – delays limited to 15 minutes
10pm to 6am – delays limited to 30 minutes
6am to 8am – delays limited to 15 minutes
Otherwise, nighttime road work at “the Narrows” will continue through August from Sunday through Thursday from 8pm to 8am with the following delays:
8pm to 10pm – delays limited to 5 minutes
10pm to 6am – delays limited to 30 minutes
6am to 8am – delays limited to 5 minutes (L. Acree)
Flags at Half-Staff
Flags will be flown at half-staff on Monday, August 14, 2017, through sunset on Wednesday, August 16, 2017, at Department of the Interior facilities in honor and memory of firefighter Brent Witham, who died in the line of duty on the Lolo Peak Fire on August 2, 2017. (K. Killian)
The Yosemite Haircare is temporarily closed. Yosemite Hospitality will share reopening information once available. (L. Cesaro)
South Fork Fire And Air Quality Update
Fire: A hazard alert message is being issued for the Wawona community. The threat is of low potential and may require evacuation in 24 hours or more. Wildfire has been discovered in the South Fork of the Merced drainage 1 mile east of Wawona. Residents and visitors are encouraged to make appropriate pre-evacuation preparations. Smoke impacts are expected to be heavy starting this evening throughout the entire community. If there is an actual need for evacuation, there will be door to door notifications by park staff.
Air Quality: Very smoky, unhealthy to hazardous conditions expected in Wawona this evening through noon Monday. Smoke from the South Fork Fire is expected to fill the canyons west of Wawona out to Mariposa as well as affecting visibility along Hwy 140 from El Portal to Mariposa and Wawona Road south to Fish Camp.
Recommendations: Close windows ASAP to keep the bulk of the smoke out. Visit friends/relatives to the north or south if you have that option.
Other areas likely to be impacted include passes to the ENE, (Tioga Pass and down to the Mono Lake area). (J. Trust)
Yosemite Fire Update Aug 14, 2017
Size: 400 acres / Elevation 5,100' / Location 37 32.258 x -119 35.902 1 mile east of Wawona
South Fork of the Merced drainage, south of Wawona Dome. Air resources on scene and suppression operations are in progress. Rangers have closed Chilnualna Falls trail at the bottom. Fire is visible from the Swinging Bridge area. Rangers have begun door to door notifications for pre-evacuation.
Multiple resources from Yosemite and adjacent communities. Joint effort with all park staff and divisions.
Size: 1200 acres / Elevation: 7,272’ / 80% of the perimeter is active burning in Red fir.
Fire burned actively today with good smoke dispersal to the north-northeast. Trails are closed from Bridalveil Creek Campground to the second junction of Ostrander trail. Additionally from Westfall Meadow to the Deer Camp junction, along Alder Creek remains closed. Signs are placed at the junctions by crews. Day-use hikers are being escorted by fire personnel at the trail near Bridalveil Creek.
Engine 32 is on the Bridalveil side of the fire and there is a new spot that is going to add .6 miles of additional trail closure. Both the trail closure and the spot are west of the creek proper. The continuation will be from the current closure to the next intersection to the south.
The crews on scene will continue to monitor fire behavior and smoke dispersion as well as assessing potential holding areas around the fire planning area.
Size: 3 acres/ Elevation: 7980’ / Fuels: Red fir and Lodge Pole. Monitored.
2-3 foot flame lengths with flames visible from Subdome. Fire is triangular shaped and active flames on both flanks. Moving uphill - made 300 foot upslope progress. Smoke is more visible than in days past, rising 300-400 ft agl and dispersing south and east.
Size: 0.16 acres / Size: less than one quarter acre / Fuels: Red fir
(.04 acre growth) 60% active perimeter with main activity to east and northeast. Creeping and smoldering in duff, litter and heavies with moderate spread potential and 2-3 ft flame lengths in the heavies. Continues to back downhill with increase in smoke production this afternoon. Smoke rising 300 ft agl and dispersing north. Continue to monitor activity and spread potential.
E-1130 monitoring this fire. Spike camping at the top of campground road.
Size: .039 acres / Elevation: 8980' / Fuels: Timber litter
20% active perimeter. Engine 1130 will be monitoring this fire. Moderate creeping and smoldering. Minimal smoke production. Will continue to monitor fire behavior. Minimal growth on east aspect
Size: 1.32 acres/ Elevation: 8154’ / Fuels: Lodge Pole and Red fir. Engine 1130 monitoring.
Update from Engine 1130: The Porcupine Fire has a 1/3 acre growth, bringing the new total to 1.32 acres. The fire has a 50% active perimeter, with primary movement to the west. Fire behavior is smoldering and creeping in duff and litter within 100 and 1000 hour fuel with single tree torching. The smoke is dissipating at 700 foot AGL to the Northeast, most active at 1400 with low RH for the day at 12%. The current plan is to monitor behavior and spread potential.
Size: less than one quarter acre/ Elevation: 10,043' / Fuels: White Bark
Unstaffed. (S. Saunders)
CONSTRUCTION & TRAFFIC DELAYS
Forestry Work Along Wawona Road
Forestry will be cutting tree hazards along the Wawona Road south of Strawberry Creek this week. Anticipate periodic traffic holds up to 15 minutes (or more, for difficult trees). Please watch out for workers, equipment, and debris along the roadway. (B. Mattos)
Wawona Road Delays
Between the Wawona Tunnel and the Rostrum: There will continue to be up to 30 minute delays Sunday night thru Thursday night, 8PM to 6AM with single lane traffic control as crews complete storm damage repairs along Wawona Road.
Near South Entrance: There will continue to be up to 30-minute daytime delays Monday thru Friday 6AM to 8PM with single lane traffic control and pilot cars. There will be extensive truck traffic over the next few weeks as crews are preparing the final subgrade for paving. (M. Pieper)